# Harnessing Diffusion-Generated Synthetic Images for Fair Image Classification

### v1

@@ 5 Conclusion
In this work, we investigated Stable Diffusion and various finetuning approaches, including Dreambooth and LoRA, to enhance fairness in image classification by generating representative images for each training group. We also explored Clustered Dreambooth, which addresses intra-group diversity by clustering images within each group and training separate Dreambooth models per cluster, preventing a single model from being overwhelmed by excessive variation. Using these approaches, we generated group-balanced images, pretrained a classifier on them, and finetuned it on real data. Experiments on three fairness benchmarks demonstrated that diffusion-based finetuning, particularly Clustered Dreambooth, consistently outperforms vanilla SD and FFR (qraitem2023fake), achieving comparable or superior results to SOTA debiasing methods like GroupDRO—especially as dataset biases become more severe.
@@ Appendix A Appendix
We organize this appendix as follows. We begin with details on the loss functions used to train our classification pipeline and the data split for the UTKFace dataset (zhang2017age). Next, we describe the prompts used for generating images from different diffusion model variants. We then present key ablations on our design choices, where we demonstrate the importance of the CLIP-Score introduced in eq. 3.2, analyze various strategies for the classification finetuning stage, and examine the impact of selecting different proportions of synthetic images for pretraining. Additionally, we discuss the roles of CE and SupCon losses in the classification pipeline. We conclude with an analysis of hyperparameter effects during pretraining with group-balanced synthetic data, an evaluation of representation similarity between Clustered Dreambooth images and real data, a report on worst-group classification performance after the pretraining stage, and qualitative examples of images generated by different methods explored in this work.
@@ A.1 Loss Function for the Two-Stage Pipeline
The SupCon loss encourages the model to learn more discriminative features by promoting greater separation between samples from different classes. We consider a mini-batch of size $\mathcal{B}$ of features denoted as $\{\mathbf{e(x_{1})},\mathbf{e(x_{2})},\cdots,\mathbf{e(x_{\mathcal{B}})}\}$ and corresponding class labels as $\{y_{1},y_{2},\dots,y_{\mathcal{B}}\}$ ($e$ is the feature encoder). Let us consider the current sample with index $j$, and the set of positive examples from the mini-batch by $P_{j}:\{i\in\mathcal{B}\quad s.t.\quad y_{i}=y_{j}\}$. Similarly, the set of negative examples is denoted by $N_{j}:\{i\in\mathcal{B}\quad s.t.\quad y_{i}\neq y_{j}\}$. The SupCon loss (khosla2020supervised) for a image $x_{j}$ is defined as,
@@ A.2 UTKFace Splits
Following previous works (qraitem2023bias; qraitem2023fake), we deliberately introduce biases into the UTKFace dataset (zhang2017age), wherein we use Gender as the target attribute and Age as the bias attribute with $90\%$ bias ratio. Particularly, we ensure that the number of images in the training set for the different groups are: $103$, $934$, $5730$ and $636$ for male adults, male children, female adults and female children respectively. To binarize the Age attribute, ages $\leq 10$ are considered as children, and those $\geq 20$ are considered adults (qraitem2023bias; qraitem2023fake). Thus, the females are biased towards older age, whereas the males are biased towards the children.
@@ A.3 Prompts used for Generation
Prompts for Vanilla SD (ramesh2022hierarchical). For Waterbirds, the prompt used is “photo of a {class-label} on {bias-label}.”, where class-label can be landbird or waterbird, and bias-label can be land or water. For UTKFace, the prompt used is “photo of a {class-label} {bias-label}.”, where class-label can be female or male, bias-label can be child or adult. Finally, for CelebA, we use the following template: “photo of a {bias-label} person with blond hair” to generate males and females (i.e., the bias-labels) with blond hair. For the non-blond class, we prompt the model with “photo of a {bias-label} person”, whereas we use a negative prompt having ‘blond hair’, to force the model to generate non-blond males and females. As the vanilla SD is not finetuned on our training sets, we use the same set of prompts for the bias ratio $0.999$ case.
Prompts for Finetuned Diffusion Models. For LoRA-finetuning, we use prompts of the format ‘Photo of a {class-label} on {bias-label}’ for Waterbirds (class-label $\in$ {waterbird, landbird}, bias-label $\in$ {water, land}), and ‘Photo of a {class-label} person who is a {bias-label}’ for UTKFace (class-label $\in$ {male, female}, bias-label $\in$ {adult, child}). For CelebA, the prompts used are: ‘Photo of a non-blond {bias-label} person’ for the class Non-Blond, and ‘Photo of a {bias-label} person with blond hair’ for the class Blond, where bias-label $\in$ {male, female}. We use the same set of prompts for Dreambooth (ruiz2022dreambooth) (i.e., single model per group) and Clustered Dreambooth. Recall that each of these methods learn specific tokens to represent the groups (or clusters in the groups in case of Clustered Dreambooth). We denote the learnt tokens by ‘[V]’. Likewise, for Waterbirds, we use the prompt “photo of a [V] bird”, where ‘[V]’ represents the learnt tokens by the model trained on the specific group or cluster. For UTKFace and CelebA, we find that providing the class-label in the prompt generates more accurate images. Hence, for UTKFace, the prompt is of the form “photo of a [V] {class-label} person”. For CelebA, the blond-class images are generated using “photo of a [V] person with blond hair”, whereas for the non-blond class, the prompt is “photo of a [V] person” with blond hair as the negative prompt.
@@ A.4 Extended Ablation Studies
Figure 6: Variations in worst group accuracies (WGA) with changing Hyperparameter values for the pretraining stage on the CelebA (celeba) and Waterbirds (sagawa2019distributionally) datasets. Apart from a drastic fall at learning rate=1e51e-5, we do not observe any significant variation in the model performances across the observe range of hyperparameters.
@@ A.5 Hyperparameters for Classification
While we do not experiment with hyperparameters during classification, using a learning rate of 1e31e-3, weight decay of 1e31e-3, batch size of $128$ and total epochs = $20$, we assess the variance of the classifier worst group performance on the pretraining stage of CelebA (celeba) and Waterbirds (sagawa2019distributionally) using the Clustered Dreambooth images. Specifically, we explore a set of learning rates : $\{1e-5,1e-4,1e-3,1e-2\}$, a range of weight decay values: $\{1e-5,1e-4,1e-3,1e-2\}$, training epochs: $\{10,20,30,40\}$, and batch sizes: $\{32,64,128,256\}$. We observe a drastic fall in accuracy scores with learning rate =1e5=1e-5 for both datasets, which is expected, considering the low value of the learning rate. While the classifier performance on Waterbirds seems more consistent across hyperparameter values than that on CelebA, we do not observe any significant variation in model performance for different hyperparameters for any of the examined datasets. We plot our findings in Fig. 6.
@@
To assess the similarities between the real and the generated distributions for each of the investigated approaches, we determine the Fréchet Inception Distance (FID) (heusel2017gans) measured between the generated images and real images of each group of CelebA. Notably, Clustered Dreambooth images outperform all competing methods by a large margin, pointing to their similarity with the training images (see Table 11).
@@ A.7 Performance of Stage 1 pretraining on Test Data
Vanilla SD (ramesh2022hierarchical) $65.4$ $79.9$ $74.6$ $80.5$
LoRA finetuning (gal2022image) $82.3$ $89.3$ $86.5$ $89.9$
Dreambooth (ruiz2022dreambooth) $85.7$ $89.5$ $89.3$ $90.1$
@@
Vanilla SD (ramesh2022hierarchical) $78.3$ $84.3$ $76.40$ $84.2$
LoRA finetuning (gal2022image) $82.1$ $87.1$ $82.3$ $87.2$
Dreambooth (ruiz2022dreambooth) $81.7$ $86.5$ $82.1$ $87.9$
@@
Vanilla SD (ramesh2022hierarchical) $68.0$ $82.3$ $62.0$ $82.3$
LoRA finetuning (gal2022image) $57.0$ $82.0$ $68.6$ $85.6$
Dreambooth (ruiz2022dreambooth) $56.2$ $80.7$ $57.9$ $80.9$
@@ A.8 Qualitative Examples
Vanilla SD failures in Waterbirds Vanilla SD models often fail to follow prompts accurately, perhaps because of inherent biases embedded in them. For example, when the model is instructed to generate ‘Photo of a waterbird on land’, many of the generated images have water in them, even when we set a negative prompt ‘water’. This problem is highlighted in Fig. 8, where more than $50\%$ of the images have water in the images. Moreover, while we agree that the generations are aesthically of high quality, the Waterbirds (sagawa2019distributionally) dataset itself is created by pasting bird images from the Caltech-UCSD Birds-200-2011 (CUB) dataset (welinder2010caltech) into images from the Place dataset (zhou2017places), which often gives the images from the original dataset an unnatural look. Thus, this leads to a domain mismatch between the training and generated images, explaining the lower worst group accuracies of Vanilla SD.

### v2

@@ 5 Conclusion
In this work, we investigated Stable Diffusion and various finetuning approaches, including Dreambooth and LoRA, to enhance fairness in image classification by generating representative images for each training group. We also explored Clustered Dreambooth, which addresses intra-group diversity by clustering images within each group and training separate Dreambooth models per cluster, preventing a single model from being overwhelmed by excessive variation. Using these approaches, we generated group-balanced images, pretrained a classifier on them, and finetuned it on real data. Experiments on three fairness benchmarks demonstrated that diffusion-based finetuning, particularly Clustered Dreambooth, consistently outperforms vanilla SD and FFR (Qraitem, Saenko, and Plummer 2023b), achieving comparable or superior results to SOTA debiasing methods like GroupDRO—especially as dataset biases become more severe.
@@ Appendix A Appendix
We organize this appendix as follows. We begin with details on the loss functions used to train our classification pipeline and the data split for the UTKFace dataset (Zhang, Song, and Qi 2017). Next, we describe the prompts used for generating images from different diffusion model variants. We then present key ablations on our design choices, where we demonstrate the importance of the CLIP-Score introduced in eq. 3.2, analyze various strategies for the classification finetuning stage, and examine the impact of selecting different proportions of synthetic images for pretraining. Additionally, we discuss the roles of CE and SupCon losses in the classification pipeline. We conclude with an analysis of hyperparameter effects during pretraining with group-balanced synthetic data, an evaluation of representation similarity between Clustered Dreambooth images and real data, a report on worst-group classification performance after the pretraining stage, and qualitative examples of images generated by different methods explored in this work.
@@ A.1 Loss Function for the Two-Stage Pipeline
The SupCon loss encourages the model to learn more discriminative features by promoting greater separation between samples from different classes. We consider a mini-batch of size $\mathcal{B}$ of features denoted as $\{\mathbf{e(x_{1})},\mathbf{e(x_{2})},\cdots,\mathbf{e(x_{\mathcal{B}})}\}$ and corresponding class labels as $\{y_{1},y_{2},\dots,y_{\mathcal{B}}\}$ ($e$ is the feature encoder). Let us consider the current sample with index $j$, and the set of positive examples from the mini-batch by $P_{j}:\{i\in\mathcal{B}\quad s.t.\quad y_{i}=y_{j}\}$. Similarly, the set of negative examples is denoted by $N_{j}:\{i\in\mathcal{B}\quad s.t.\quad y_{i}\neq y_{j}\}$. The SupCon loss (Khosla et al. 2020) for a image $x_{j}$ is defined as,
@@ A.2 UTKFace Splits
Following previous works (Qraitem, Saenko, and Plummer 2023a, b), we deliberately introduce biases into the UTKFace dataset (Zhang, Song, and Qi 2017), wherein we use Gender as the target attribute and Age as the bias attribute with $90\%$ bias ratio. Particularly, we ensure that the number of images in the training set for the different groups are: $103$, $934$, $5730$ and $636$ for male adults, male children, female adults and female children respectively. To binarize the Age attribute, ages $\leq 10$ are considered as children, and those $\geq 20$ are considered adults (Qraitem, Saenko, and Plummer 2023a, b). Thus, the females are biased towards older age, whereas the males are biased towards the children.
@@ A.3 Prompts used for Generation
Prompts for Vanilla SD (Ramesh et al. 2022). For Waterbirds, the prompt used is “photo of a {class-label} on {bias-label}.”, where class-label can be landbird or waterbird, and bias-label can be land or water. For UTKFace, the prompt used is “photo of a {class-label} {bias-label}.”, where class-label can be female or male, bias-label can be child or adult. Finally, for CelebA, we use the following template: “photo of a {bias-label} person with blond hair” to generate males and females (i.e., the bias-labels) with blond hair. For the non-blond class, we prompt the model with “photo of a {bias-label} person”, whereas we use a negative prompt having ‘blond hair’, to force the model to generate non-blond males and females. As the vanilla SD is not finetuned on our training sets, we use the same set of prompts for the bias ratio $0.999$ case.
Prompts for Finetuned Diffusion Models. For LoRA-finetuning, we use prompts of the format ‘Photo of a {class-label} on {bias-label}’ for Waterbirds (class-label $\in$ {waterbird, landbird}, bias-label $\in$ {water, land}), and ‘Photo of a {class-label} person who is a {bias-label}’ for UTKFace (class-label $\in$ {male, female}, bias-label $\in$ {adult, child}). For CelebA, the prompts used are: ‘Photo of a non-blond {bias-label} person’ for the class Non-Blond, and ‘Photo of a {bias-label} person with blond hair’ for the class Blond, where bias-label $\in$ {male, female}. We use the same set of prompts for Dreambooth (Ruiz et al. 2022) (i.e., single model per group) and Clustered Dreambooth. Recall that each of these methods learn specific tokens to represent the groups (or clusters in the groups in case of Clustered Dreambooth). We denote the learnt tokens by ‘[V]’. Likewise, for Waterbirds, we use the prompt “photo of a [V] bird”, where ‘[V]’ represents the learnt tokens by the model trained on the specific group or cluster. For UTKFace and CelebA, we find that providing the class-label in the prompt generates more accurate images. Hence, for UTKFace, the prompt is of the form “photo of a [V] {class-label} person”. For CelebA, the blond-class images are generated using “photo of a [V] person with blond hair”, whereas for the non-blond class, the prompt is “photo of a [V] person” with blond hair as the negative prompt.
@@ A.4 Extended Ablation Studies
Figure 6: Variations in worst group accuracies (WGA) with changing Hyperparameter values for the pretraining stage on the CelebA (Celeba ) and Waterbirds (Sagawa et al. 2019) datasets. Apart from a drastic fall at learning rate=1e51e-5, we do not observe any significant variation in the model performances across the observe range of hyperparameters.
@@ A.5 Hyperparameters for Classification
While we do not experiment with hyperparameters during classification, using a learning rate of 1e31e-3, weight decay of 1e31e-3, batch size of $128$ and total epochs = $20$, we assess the variance of the classifier worst group performance on the pretraining stage of CelebA (Celeba ) and Waterbirds (Sagawa et al. 2019) using the Clustered Dreambooth images. Specifically, we explore a set of learning rates : $\{1e-5,1e-4,1e-3,1e-2\}$, a range of weight decay values: $\{1e-5,1e-4,1e-3,1e-2\}$, training epochs: $\{10,20,30,40\}$, and batch sizes: $\{32,64,128,256\}$. We observe a drastic fall in accuracy scores with learning rate =1e5=1e-5 for both datasets, which is expected, considering the low value of the learning rate. While the classifier performance on Waterbirds seems more consistent across hyperparameter values than that on CelebA, we do not observe any significant variation in model performance for different hyperparameters for any of the examined datasets. We plot our findings in Fig. 6.
@@
To assess the similarities between the real and the generated distributions for each of the investigated approaches, we determine the Fréchet Inception Distance (FID) (Heusel et al. 2017) measured between the generated images and real images of each group of CelebA. Notably, Clustered Dreambooth images outperform all competing methods by a large margin, pointing to their similarity with the training images (see Table 11).
@@ A.7 Performance of Stage 1 pretraining on Test Data
Vanilla SD (Ramesh et al. 2022) $65.4$ $79.9$ $74.6$ $80.5$
LoRA finetuning (Gal et al. 2022) $82.3$ $89.3$ $86.5$ $89.9$
Dreambooth (Ruiz et al. 2022) $85.7$ $89.5$ $89.3$ $90.1$
@@
Vanilla SD (Ramesh et al. 2022) $78.3$ $84.3$ $76.40$ $84.2$
LoRA finetuning (Gal et al. 2022) $82.1$ $87.1$ $82.3$ $87.2$
Dreambooth (Ruiz et al. 2022) $81.7$ $86.5$ $82.1$ $87.9$
@@
Vanilla SD (Ramesh et al. 2022) $68.0$ $82.3$ $62.0$ $82.3$
LoRA finetuning (Gal et al. 2022) $57.0$ $82.0$ $68.6$ $85.6$
Dreambooth (Ruiz et al. 2022) $56.2$ $80.7$ $57.9$ $80.9$
@@ A.8 Qualitative Examples
Vanilla SD failures in Waterbirds Vanilla SD models often fail to follow prompts accurately, perhaps because of inherent biases embedded in them. For example, when the model is instructed to generate ‘Photo of a waterbird on land’, many of the generated images have water in them, even when we set a negative prompt ‘water’. This problem is highlighted in Fig. 8, where more than $50\%$ of the images have water in the images. Moreover, while we agree that the generations are aesthically of high quality, the Waterbirds (Sagawa et al. 2019) dataset itself is created by pasting bird images from the Caltech-UCSD Birds-200-2011 (CUB) dataset (Welinder et al. 2010) into images from the Place dataset (Zhou et al. 2017), which often gives the images from the original dataset an unnatural look. Thus, this leads to a domain mismatch between the training and generated images, explaining the lower worst group accuracies of Vanilla SD.